OpenStudy (anonymous):

solve the equation step by step. list each step, along with the property you used during that step. -5*1*11*4

OpenStudy (anonymous):

-5*1=-5 -5*11=-55 -55*4=-220

OpenStudy (anonymous):

can you help out by tellin me the property u used

OpenStudy (anonymous):

all steps are multiplication so you can start the step which you want